Vector A is 3.0 units in length and points along the positive x-axis; vector B is 4.0 units in length and points along a direction 150 from the positive x-axis. What is the magnitude of the resultant when vectors A and B are added?


1
Expert's answer
2021-10-13T09:06:32-0400

Find the x- and y-components of both vectors, then find the x- and y-components of the resultant, then find the magnitude of the resultant itself:


"A_x=3,\\
A_y=0.\\
B_x=B\cos\theta=4\cos150°=-3.46,\\
B_y=B\sin\theta=4\sin150°=2.\\\space\\
R_x=A_x+B_x,\\
R_y=A_y+B_y,\\
R=\sqrt{R_x^2+R_y^2}=2.05."